# EEG-Language Modeling for Pathology Detection

### v3

@@ B.3 Language Encoder
We compare three pretrained language models in their ability to perform zero-shot pathology detection following EEG-language pretraining (Table 9). We find that MedCPT performs best [40], which is trained using contrastive learning with 255 million user click logs from PubMed.
@@ B.5 EEG-Only Pretraining
Bootstrap-Your-Own-Latent. BYOL relies on two encoder models: an online and a target network [41]. During pretraining, the online network is trained to predict the target model’s output. Meanwhile, the weights of the target network are updated using a moving average of the weights of the online network, which has been empirically shown to prevent collapse of the latent space. For alignment, $\ell_{2}$ normalization is applied to the EEG embeddings $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ and the mean square distance is minimized. We adopt the recommended parameter value for the exponential moving average [41]. The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 32.
@@
Variance-Invariance-Covariance Regularization. VICReg allows for the use of a single encoder model and prevents collapse by applying two explicit regularization terms to each of the embedding batches $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ [42]. The ’variance’ term maintains the standard deviation (computed batch-wise) of every embedding dimension above a threshold, thereby avoiding a trivial solution. In addition, latent collapse is avoided through the ’covariance’ term which decorrelates pairs of embedding dimensions. The method minimizes the mean square distance between $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$. Hyperparameters are set to their recommended values [42]. The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 256.
@@
Contrast with the World Representation. ContraWR was proposed to improve augmentation-based SSL for EEG [9]. The method, which is contrastive in nature, maximizes similarity between $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ while preventing collapse by minimizing similarity with ’negative’ samples. ContraWR forms a negative representation by aggregating across all negative batch elements, aiming to compensate for the low signal-to-noise of EEG data by creating a more reliable negative contrast. It relies on a triplet loss based on Info-NCE [53]. We also here set the hyperparameters to the values recommended by the authors [9]. The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 32.
@@
Relative Positioning. Pairs of EEG crops are sampled and assigned binary labels based on their temporal proximity [10]. Crops close in time are labeled positive, while those far apart are labeled negative. We use the same EEG encoder as for all other methods to create representations and use the suggested contrastive module to compute the element-wise absolute difference between representations. A logistic regression model then predicts the label. The method is trained using binary logistic loss. For all methods by [10], we use the hyperparameters reported to work best on TUAB, including between-subject sampling of EEG crops.
@@
Temporal Shuffling. An extension of Relative Positioning by sampling triplets of EEG crops. The task is to determine whether the crops are in temporal order or shuffled [10]. The contrastive module concatenates absolute differences between representations. As with Relative Positioning, a logistic regression model is used for prediction, and the method is trained end-to-end using binary logistic loss.
@@
Contrastive Predictive Coding. This method uses an autoregressive encoder to summarize a sequence of EEG crops into a context vector [10]. The task is to predict which future crop actually follows the context, among negative samples. A bilinear model is used for prediction at each future step. The method is trained end-to-end using the InfoNCE loss.
@@ B.5.1 Data augmentations
For EEG-only pretraining, we adapt the data augmentations proposed by [8], which were found to perform well on the TUAB dataset. For a given EEG crop, we apply the same augmentation to each channel. Parameters are sampled independently for each EEG crop and uniformly from the ranges displayed in table 10. Augmentations are visualized for a single EEG channel in figure 9.
@@
Table 10: Data augmentation parameter ranges; adapted from [8].
@@ C.1 EEG preprocessing
EEG data received minimal preprocessing (using MNE [54]). First, the initial 10 seconds were removed to reduce the impact of set-up artefacts. Afterwards, a bandpass filter of 0.1-49 Hz was applied and all recordings were resampled to 100 Hz. To reduce the impact of signal artefacts, all EEG signals had their amplitude clipped to $\pm$ 800 $\mu$V. As a large majority of recordings used an average-reference (AR) or linked-ear reference (LE), we only used these recordings and standardized them via transformation to the 20-channel Temporal Central Parasagittal (TCP) montage.
@@ C.2 Subsampling
TUEG contains considerably more abnormal than normal EEGs. As vision-language models have been shown to be sensitive to imbalanced classes [55], we subsample the data to create approximately equal class representation. We rely on the LLM summaries of reports, which were more consistent in their phrasing regarding the normal or abnormal status. This allowed for a more reliable classification using regular expressions. All reports for which no clear classification was made were omitted. 5015 reports in the potential training set were classified as normal, which were associated with 7526 EEG recordings. For our ‘pretrain’ data subset, we subsampled the abnormal EEGs to match the amount of normal EEG recordings. This resulted in 7526 abnormal EEG recordings, with 6770 reports. Although only a minor subset of these preliminary classifications was manually verified, it is important to note that this process was solely to alleviate severe class imbalance and was not used for further analysis.
@@
To test for retrieval performance, we supplemented the TUAB test set with data from the TUH EEG Epilepsy Corpus [56] in an attempt to create a larger, roughly balanced evaluation set of those with and without pathology. For this, we only selected the first recording of a subject so that no multiple files from the same subject were present. Additionally, we only included reports which had at least one heading from each text cluster to allow for a fair comparison.
@@ Appendix D Classification
For linear evaluation, we train logistic linear regression models using 10-fold cross validation for each pretrained model using sklearn [57]. We perform grid-search over 45 logarithmically-spaced values for L2 regularization between $10^{-6}$ and $10^{5}$ via a validation set.
@@
For zero-shot pathology detection, we perform an ensemble over 21 binary prompts, listed in Table 11. Prompt ensembling was shown to improve performance [12], but we employ it here also as the limited data is likely to lead to less stable representations, which may lead to sensitivity to phrasing. To inspect whether results are sensitive to changes to the prompt set, we perform a post-hoc analysis using the held-out test set that iteratively leaves one prompt out of the ensemble (Figure 10). We observe that results are consistent across such reduced prompt sets, except for the ELMl model trained on the clinical history or interpretation clusters, although neither model reaches competitive performance. This set was only initially verified on the training set to enable model- and parameter-comparisons using zero-shot performance. Tuning is likely to enable further performance improvements, although the flexibility of the zero-shot approach may introduce severe risk of overfitting on the TUEG dataset.
@@ E.1 EEG embeddings of pathology
We provide t-SNE (complexity=40, [58]) visualisations of the averaged EEG embeddings per subject after pretraining. These are post-hoc plots for which we use models trained on the entire pretraining subset and display embeddings of hold-out TUAB patients. ELMe,l and ELM-MIL show the clearest visual separation between abnormal and normal EEGs, which is in line with the linear probing results.
@@ E.2 Within-subject EEG embeddings
E.3 Report and content segmentation

### v4

@@ B.3 Language Encoder
We compare three pretrained language models in their ability to perform zero-shot pathology detection following EEG-language pretraining (Table 12). We find that MedCPT performs best (Jin et al., 2023), which is trained using contrastive learning with 255 million user click logs from PubMed.
@@ B.5 EEG-Only Pretraining
Bootstrap-Your-Own-Latent. BYOL relies on two encoder models: an online and a target network (Grill et al., 2020). During pretraining, the online network is trained to predict the target model’s output. Meanwhile, the weights of the target network are updated using a moving average of the weights of the online network, which has been empirically shown to prevent collapse of the latent space. For alignment, $\ell_{2}$ normalization is applied to the EEG embeddings $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ and the mean square distance is minimized. We adopt the recommended parameter value for the exponential moving average (Grill et al., 2020). The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 32.
@@
Variance-Invariance-Covariance Regularization. VICReg allows for the use of a single encoder model and prevents collapse by applying two explicit regularization terms to each of the embedding batches $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ (Bardes et al., 2021). The ’variance’ term maintains the standard deviation (computed batch-wise) of every embedding dimension above a threshold, thereby avoiding a trivial solution. In addition, latent collapse is avoided through the ’covariance’ term which decorrelates pairs of embedding dimensions. The method minimizes the mean square distance between $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$. Hyperparameters are set to their recommended values (Bardes et al., 2021). The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 256.
@@
Contrast with the World Representation. ContraWR was proposed to improve augmentation-based SSL for EEG (Yang et al., 2021). The method, which is contrastive in nature, maximizes similarity between $\{\textbf{h}_{e}^{\prime},\textbf{h}_{e}^{\prime\prime}\}$ while preventing collapse by minimizing similarity with ’negative’ samples. ContraWR forms a negative representation by aggregating across all negative batch elements, aiming to compensate for the low signal-to-noise of EEG data by creating a more reliable negative contrast. It relies on a triplet loss based on Info-NCE (Gutmann & Hyvärinen, 2010). We also here set the hyperparameters to the values recommended by the authors (Yang et al., 2021). The projection head is a 2-layer non-linear MLP with a hidden dimension of width 256 and an output dimension of 32.
@@
Relative Positioning. Pairs of EEG crops are sampled and assigned binary labels based on their temporal proximity (Banville et al., 2021). Crops close in time are labeled positive, while those far apart are labeled negative. We use the same EEG encoder as for all other methods to create representations and use the suggested contrastive module to compute the element-wise absolute difference between representations. A logistic regression model then predicts the label. The method is trained using binary logistic loss. For all methods by (Banville et al., 2021), we use the hyperparameters reported to work best on TUAB, including between-subject sampling of EEG crops.
@@
Temporal Shuffling. An extension of Relative Positioning by sampling triplets of EEG crops. The task is to determine whether the crops are in temporal order or shuffled (Banville et al., 2021). The contrastive module concatenates absolute differences between representations. As with Relative Positioning, a logistic regression model is used for prediction, and the method is trained end-to-end using binary logistic loss.
@@
Contrastive Predictive Coding. This method uses an autoregressive encoder to summarize a sequence of EEG crops into a context vector (Banville et al., 2021). The task is to predict which future crop actually follows the context, among negative samples. A bilinear model is used for prediction at each future step. The method is trained end-to-end using the InfoNCE loss.
@@ B.5.1 Data augmentations
For EEG-only pretraining, we adapt the data augmentations proposed by Mohsenvand et al. (2020), which were found to perform well on the TUAB dataset. For a given EEG crop, we apply the same augmentation to each channel. Parameters are sampled independently for each EEG crop and uniformly from the ranges displayed in table 13. Augmentations are visualized for a single EEG channel in figure 10.
@@
Table 13: Data augmentation parameter ranges; adapted from Mohsenvand et al. (2020).
@@ C.1 EEG preprocessing
[…]
@@ C.2 Subsampling
TUEG contains considerably more abnormal than normal EEGs. As vision-language models have been shown to be sensitive to imbalanced classes (Wang et al., 2024), we subsample the data to create approximately equal class representation. We rely on the LLM summaries of reports, which were more consistent in their phrasing regarding the normal or abnormal status. This allowed for a more reliable classification using regular expressions. All reports for which no clear classification was made were omitted. 5015 reports in the potential training set were classified as normal, which were associated with 7526 EEG recordings. For our ‘pretrain’ data subset, we subsampled the abnormal EEGs to match the amount of normal EEG recordings. This resulted in 7526 abnormal EEG recordings, with 6770 reports. Although only a minor subset of these preliminary classifications was manually verified, it is important to note that this process was solely to alleviate severe class imbalance and was not used for further analysis.
@@
To test for retrieval performance, we supplemented the TUAB test set with data from the TUH EEG Epilepsy Corpus (Veloso et al., 2017) in an attempt to create a larger, roughly balanced evaluation set of those with and without pathology. For this, we only selected the first recording of a subject so that no multiple files from the same subject were present. Additionally, we only included reports which had at least one heading from each text cluster to allow for a fair comparison.
@@ Appendix D Classification
For linear evaluation, we train logistic linear regression models using 10-fold cross validation for each pretrained model using sklearn (Pedregosa et al., 2011). We perform grid-search over 45 logarithmically-spaced values for L2 regularization between $10^{-6}$ and $10^{5}$ via a validation set.
@@
For zero-shot pathology detection, we perform an ensemble over 21 binary prompts, listed in Table 14. Prompt ensembling was shown to improve performance (Radford et al., 2021), but we employ it here also as the limited data is likely to lead to less stable representations, which may lead to sensitivity to phrasing. To inspect whether results are sensitive to changes to the prompt set, we perform a post-hoc analysis using the held-out test set that iteratively leaves one prompt out of the ensemble (Figure 11). We observe that results are consistent across such reduced prompt sets, except for the ELMl model trained on the clinical history or interpretation clusters, although neither model reaches competitive performance. This set was only initially verified on the training set to enable model- and parameter-comparisons using zero-shot performance. Tuning is likely to enable further performance improvements, although the flexibility of the zero-shot approach may introduce severe risk of overfitting on the TUEG dataset.
@@ E.1 EEG embeddings of pathology
We provide t-SNE (complexity=40, (Van der Maaten & Hinton, 2008)) visualisations of the averaged EEG embeddings per subject after pretraining. These are post-hoc plots for which we use models trained on the entire pretraining subset and display embeddings of hold-out TUAB patients. ELMe,l and ELM-MIL show the clearest visual separation between abnormal and normal EEGs, which is in line with the linear probing results.
@@ E.2 Within-subject EEG embeddings
E.3 Alignment visualizations
Figure 14: Additional alignment visualizations as in Figure 4.
Figure 15: Additional alignment visualizations as in Figure 4.
Figure 16: Additional alignment visualizations as in Figure 4, but for two observations which were not well aligned. Both feature a narrow range of similarity values, indicating a lack of temporal localization. Whereas mild diffuse slowing may have not been pronounced enough, left frontotemporal spikes are very rarely described in the dataset.
Figure 17: Additional alignment visualizations as in Figure 4, but for two recordings which were classified as normal by the physician.
E.4 Report and content segmentation